# Impact of flavour coupling on $SO(10)$-inspired leptogenesis

### v1

@@ 3.2.5 Majorana phases
In the right panel of Fig. 6 we show the constraints in the plane $\rho$ versus $\sigma$. This plot is quite interesting because it clearly shows how the new solutions, both muonic and electronic, stem from the tauonic $C\!P$ asymmetry. The regions they fill are completely disconnected from the primary muonic solutions that are present also in the absence of flavour coupling and for which the asymmetry is proportional to the muon $C\!P$ asymmetry. One can also notice how the new solutions extend to higher values of $\sigma$ for $\rho\sim\pi$. The same conclusions can be also drawn looking at the right panel of Fig. 7 where we show the constraints in the plane $\sigma$ versus $m_{1}$.
@@ 3.2.6 $0\nu\beta\beta$ effective neutrino mass
Finally, one can see from the right panel of Fig. 8 that flavour coupling does not really change the constraints in the plane $m_{ee}$ versus $m_{1}$. The new solutions are confined within the region that was anyway accessible also without accounting for flavour coupling.
@@ 4 Strong thermal $SO(10)$-inspired leptogenesis
In this section, we finally discuss ST-SO10INLEP [30]. It is intriguing that the rather special conditions that are needed to realise strong thermal leptogenesis [31, 32], are naturally satisfied by a subset of the solutions realising successful SO10INLEP. In this Section we first re-derive, and update, the predictions on low energy neutrino parameters from ST-SO10INLEP neglecting flavour coupling, and then we discuss the impact of flavour coupling.
@@ 4.1 Neglecting flavour coupling
Let us assume that some external mechanism has created an initial pre-existing B-L asymmetry $N_{B-L}^{\rm p,i}$ prior to leptogenesis and prior to any washout from RH neutrinos. We also assume that the $N_{3}$ mass is larger than the reheat temperature, so that the $N_{3}$-washout is absent. [^11] This assumption is conservative but ultimately it is irrelevant whether $N_{3}$-washout is included or not, the predictions on the low energy neutrino parameters change just slightly [30]. Because of the RH neutrino inverse processes, the pre-existing asymmetry will undergo a dynamical evolution that has to take into account also the different flavour regimes. Finally, the relic value of the pre-existing asymmetry can be also expressed as the sum of three contributions from each flavour i.e., $N_{B-L}^{\rm p,f}=\sum_{\alpha}\,N_{\Delta_{\alpha}}^{\rm p,f}$. Neglecting flavour coupling, the expressions for the $N_{\Delta_{\alpha}}^{\rm p,f}$ are given by [30, 40]
@@
In this expression the quantities $p^{0}_{{\rm p}\tau}$ and $p^{0}_{{\rm p}\tau_{2}^{\bot}}$ denote the fractions of the initial pre-existing asymmetry in the tauon flavour and in the flavour $\tau_{2}^{\bot}$, where, as already defined, $\tau_{2}^{\bot}$ is the electron and muon flavour superposition component in the leptons produced by $N_{2}$-decays (or equivalently the flavour component that is washed-out in the inverse processes producing $N_{2}$). [^12] Notice that in general $p^{0}_{{\rm p}\tau}+p^{0}_{{\rm p}\tau_{2}^{\bot}}\neq 1$ since in general the projection of the pre-existing asymmetry flavour direction on the $\tau^{\bot}$ plane, is not parallel to $\tau^{\bot}_{2}$. We also introduced the two quantities $p^{0}_{\alpha\tau_{2}^{\bot}}\equiv K_{2\alpha}/(K_{2e}+K_{2\mu})\;(\alpha=e,\mu)$. These give the fractions of the $\alpha$ flavour in the $\tau_{2}^{\bot}$ component of the lepton state $|\ell_{2}\rangle$ produced by $N_{2}$-decays. The terms $\Delta p_{{\rm p}\tau},\Delta p_{{\rm p}e}$ and $\Delta p_{{\rm p}\mu}$ are the so-called phantom terms and such that $\Delta p_{{\rm p}e}+\Delta p_{{\rm p}\mu}+\Delta p_{{\rm p}\tau}=0$. They are phantom terms taking into account the possibility that the mechanism that created the pre-existing asymmetry produced lepton states that are not exactly the $C\!P$ conjugated of the anti-lepton states. It can be crossed-checked that in Eq. (49), if one switches-off the washout terms, then the sum of the three final flavoured pre-existing asymmetries is equal to the total initial pre-existing asymmetry, as it should be.
@@
The strong thermal leptogenesis condition requires that the relic value of the pre-existing asymmetry is negligible with respect to the final value of the asymmetry produced by $N_{2}$-decays. For definitiness, we will impose a condition $N_{B-L}^{\rm p,f}<0.1\,N_{B-L}^{\rm lep,f}$.
@@
The only possibility to have successful strong thermal leptogenesis is just the $N_{2}$ tauon-dominated leptogenesis scenario [31] that is also representing the bulk of solutions in SO10INLEP, as we discussed. This is quite an intriguing coincidence. The reason why only tauon-dominates solutions can realise strong thermal leptogenesis can be understood directly from Eq. (49). One can indeed have in this case $K_{2\tau},K_{1\mu},K_{1e}\gg 1$ and $K_{1\tau}\lesssim 1$ in a way to wash out at $T\sim M_{2}$ the tauonic component of the pre-existing asymmetry and at $T\sim M_{1}$ also the electronic and muonic components. At the same time, having $K_{1\tau}\lesssim 1$, the tauonic asymmetry produced from leptogenesis is the only that survives the $N_{1}$-washout.
@@
There is only a very special caveat [40]. If one imposes $(1-p^{0}_{\mu\tau_{2}^{\bot}})=K_{2e}/(K_{2e}+K_{2\mu})\lesssim 0.1N_{B-L}^{\rm lep,f}/N_{B-L}^{\rm p,i}\lll 1$ and $K_{1\mu}\lesssim 1$, instead of $K_{1\mu}\gg 1$, one can have a strong thermal muon-dominated solution. We noticed that $K_{2e}/(K_{2e}+K_{2\mu})\sim\theta_{\rm c}^{2}\sim 0.05$ is indeed small but in this case, for a meaningfully large value of pre-existing asymmetry $\sim 10^{-3}$, as we will assume, one needs this ratio at least four orders of magnitude smaller. These are incredibly special solutions, unless they can be justified within a particular model. In [40] these solutions were found with a tiny $\sim 10^{-9}$ rate. They have to be considered then more like oddities rather than real ST-SO10INLEP solutions.
@@
By imposing a strong thermal leptogenesis condition in addition to successful SO10INLEP, one singles out a special subset of solutions within those we obtained just imposing successful SO10INLEP. We have assumed as benchmark value $N^{\rm p,i}_{B-L}=10^{-3}$. In the scatter plots, in Fig. 1 and Figs. 3–9, these solutions are denoted by blue points. The success rate for these solutions in our scan is $10^{-6}$. The results fully confirm previous results [30, 12, 33, 40]. It is still useful to notice some important aspects:
@@
  • From Fig. 6 one can see how one the requested values of the Majorana phases for ST-SO10INLEP become so constrained that ST-SO10INLEP would be a way to determine their values.
[…]
  • The bulk of solutions realising ST-SO10INLEP are of $\tau_{A}$ type. The $\tau_{B}$ type solutions have the problem that $K_{1\mu}$ tends to be too small to washout the muonic component of the pre-existing asymmetry. For example for $V_{L}=I$ one finds $K_{1\mu}\lesssim 4$ [12] and in this case they would be absolutely unable to yield strong thermal solutions. However, turning on angles in $V_{L}$ values as large as $K_{1\mu}\sim 10$ are possible and for this reason some $\tau_{B}$ strong thermal solutions do exist, though they are only $\sim 1\%$ of the $\tau_{B}$ solutions.
@@
Figure 10: ST-SO10INLEP allowed region in the plane $\delta$ versus $\theta_{23}$. The red line delimits the region when flavour coupling is neglected, the coloured region is the result in the case when flavour coupling is included. The colour code denotes point density: lighter blue means higher density of solutions. The thin dashed lines are the $1\sigma,2\sigma$ and $3\sigma$ experimental constraints from [23].
The results is shown with a red line in Fig. 10. This well reproduces the result in [40]. The small differences can be ascribed to the updated experimental ranges for neutrino oscillation parameters in Eq. (6) we are adopting. The thin dashed lines are the the $1\sigma,2\sigma$ and $3\sigma$ experimental constraints from [23]. One can see how ST-SO10INLEP is in nice agreement with crrent experimental results within $1\sigma$. The fact that currently the best fit is in the first octant releases the tension that was found with old experimental data in [40]. The best fit $\delta$ value is actually too low but this is very sensitive to the employed data sets in the analysis, mainly because of an existing tension between T2K and NOVA data. In the left panel of Fig. 11 we show
@@
plots of $\eta_{B}^{\rm lep}$ versus $m_{1}$ for the indicated values of $\theta_{23}$ and for fixed $\delta=-50^{\circ}$, corresponding to the value allowing the maximum value of $\theta_{23}$. All other parameters are fixed to values that maximise the asymmetry and are given in Table 3. One can see how for increasing values of $\theta_{23}$ the peak value of $\eta_{B}^{\rm lep}$ decreases. This is well understood since the asymmetry is approximately $\propto 1/\sin^{4}\theta_{23}$ [12, 33]. One can see how the maximum value of $\theta_{23}$ we obtain is $46^{\circ}$. This is consistent with the value obtained in [40] and it clearly shows how second octant is incompatible with ST-SO10INLEP.
@@
No F.C. $33.62^{\circ}$ $9.09^{\circ}$ See Figure $-49^{\circ}$ 0.312 0.834 $4.11^{\circ}$ $0.183^{\circ}$ $2.41^{\circ}$ 1.92 1.24 0.18
With F.C. $33.62^{\circ}$ $9.09^{\circ}$ See Figure $-49^{\circ}$ 0.312 0.834 $4.11^{\circ}$ $0.183^{\circ}$ $2.41^{\circ}$ 1.92 1.24 0.18
@@ 4.2 Including flavour coupling
We have now finally to discuss the impact of flavour coupling on ST-SO10INLEP. This is one of the main motivations of the paper. A naive, but legitimate, expectation is that flavour coupling might completely disrupt ST-SO10INLEP. The reason is that a large pre-existing asymmetry, thanks to flavour coupling, can find its way to survive into the tauon flavour where the washout is weak. As we are going to explain, this expectation is incorrect.
We have first of all to generalise the expression (49) for the relic value of the pre-existing asymmetry. If we start with an initial pre-existing asymmetry $N^{\rm p,i}_{B-L}$, that we assume to be generated at some temperature above $\sim 10^{12}\,{\rm GeV}$, below $10^{12}\,{\rm GeV}$ the tauon component of lepton states is measured by tauon charged lepton interactions and the pre-existing asymmetry will be the sum of a tauon component, given by $p^{0}_{{\rm p}\tau}\,N^{\rm p,i}_{B-L}$, and a component along a flavour direction given by the projection of the flavour composition of pre-existing leptons on the $e$-$\mu$ flavour plane given by $(1-p^{0}_{{\rm p}\tau})\,N^{\rm p,i}_{B-L}$. In addition, in general, there should be also the phantom terms that however we neglect just to shorten the notation (in any case again results are not sensitive to them). Similarly to the solution for the asymmetry generated by leptogenesis, at $T\sim M_{2}$ the $N_{2}$-washout, in the presence of flavour coupling, uncouples in a rotated flavour basis $\tau^{\prime}$ and $\tau^{\bot^{\prime}}_{2}$. The difference is that now there is no source term and, therefore, at the end of the washout processes the washout is described by exponential factors and, rotating back to the standard flavour basis, one finds
@@
where the matrix $U$ is the same we introduced in Eq. (45) when we introduced flavour coupling for the asymmetry produced by $N_{2}$-decays. It can be noticed that, for $U=I$, one recovers the result in the absence of flavour coupling and $N_{1}$-washout in Eq. (49). Also notice that the component of the pre-existing asymmetry along the flavour $\tau_{2^{\bot}}^{\bot}$, orthogonal to $\tau$ and $\tau_{2}^{\bot}$, escapes the washout at $T\sim M_{2}$ [34, 55].
@@
Using now this much more intricate expression, we have obtained ST-SO10INLEP solutions shown, again with blue points, in the right panels of Fig. 1 and Figs. 2–8. The first thing to notice is that the region survives and in fact it even slightly increases. In particular, the most noticeable change is that the lower bound on $m_{1}$ gets relaxed and from $m_{1}\gtrsim 8\,{\rm meV}$, obtained neglecting flavour coupling, we have now found $m_{1}\gtrsim 3\,{\rm meV}$. One can also see from Fig. 8 that, corresponding, the lower bound on $m_{ee}$ gets relaxed from $m_{ee}\gtrsim 9\,{\rm meV}$ to $m_{ee}\gtrsim 4\,{\rm meV}$. One can also see from Fig. 6 that the two disconnected subregions now get linked by the appearance of the new solutions.
This implies, that in fact the washout of a pre-existing asymmetry, gets even easier when flavour coupling is taken into account, contrarily to naive expectations. The reasons why that happens were already partially envisaged in [33]. Even though flavour coupling induces part of the pre-existing asymmetry to be transferred in the tauon flavour where the washout is absent ($K_{1\tau}\ll 1$) to let the asymmetry produced from leptogenesis to survive, the off-diagonal term that is responsible for the coupling is proportional to $K_{1\tau}$ itself and so intrinsically small. That means that the same conditions that are necessary to have strong thermal leptogenesis are also sufficient to protect it from the potential disruptive effect of flavour coupling. In addition flavour coupling makes in a way that the flavoured decay parameters responsible for the washout of the pre-existing electron and muon asymmetries are not the standard $K_{1e}$ and $K_{1\mu}$ but these are replaced by $K_{1e}^{\prime\prime}$ and $K_{1\mu}^{\prime\prime}$. The origin of the lower bound on $m_{1}$ was studied in detail in [32, 12] and it was due to the fact that $K_{1e}\simeq m_{ee}/m_{\star}\simeq 0.8\,m_{1}/m_{ee}$. For an initial electronic pre-existing asymmetry $\sim 10^{-4}$, the strong thermal conditions implies $K_{1e}\gtrsim 8$ that translates immediately into $m_{ee}\gtrsim 8\,{\rm meV}$ and $m_{1}\gtrsim 9\,{\rm meV}$. However, when flavour coupling is included the washout of the electron asymmetry is described by $K_{1e}^{\prime\prime}$ that is a linear combination primarily of $K_{1e}$ but with a small contribution from $K_{1\mu}\simeq(m_{\rm sol}s^{2}_{12}+m_{\rm atm}s_{13}c_{23})^{2}/m_{\star}m_{\rm sol}s^{2}_{12}\simeq 25$. This contamination from a large $K_{1\mu}$ makes possible to have smaller values of $K_{1e}$ and, therefore, of $m_{ee}$ and $m_{1}$.

### v2

@@ 3.2.5 Majorana phases
In the right panel of Fig. 6 we show the constraints in the plane $\rho$ versus $\sigma$. This plot is quite interesting because it clearly shows how the new solutions, both muonic and electronic, stem from the tauonic CPC\!P asymmetry. The regions they fill are completely disconnected from the primary muonic solutions that are present also in the absence of flavour coupling and for which the asymmetry is proportional to the muon CPC\!P asymmetry. One can also notice how the new solutions extend to higher values of $\sigma$ for $\rho\sim\pi$. The same conclusions can be also drawn looking at the right panel of Fig. 7 where we show the constraints in the plane $\sigma$ versus $m_{1}$.
@@ 3.2.6 $0\nu\beta\beta$ effective neutrino mass
Finally, one can see from the right panel of Fig. 8 that flavour coupling does not change much the constraints in the plane $m_{ee}$ versus $m_{1}$. The new solutions are confined within the region that was anyway accessible also without accounting for flavour coupling. The most remarkable difference is that the lower bound on $m_{ee}$ gets relaxed by about one order-of-magnitude but it is still exists.
@@ 4 Strong thermal $SO(10)$-inspired leptogenesis
In this section, we finally discuss ST-SO10INLEP ST . It is intriguing that the rather special conditions that are needed to realise strong thermal leptogenesis problem , sophie , are naturally satisfied by a subset of the solutions realising successful SO10INLEP. In this section we first re-derive, and update, the predictions on low energy neutrino parameters from ST-SO10INLEP neglecting flavour coupling, and then we discuss the impact of flavour coupling.
@@ 4.1 Neglecting flavour coupling
Let us assume that some external mechanism has created an initial pre-existing B-L asymmetry $N_{B-L}^{\rm p,i}$ prior to leptogenesis and prior to any washout from RH neutrinos. We also assume that the $N_{3}$ mass is larger than the reheat temperature, so that the $N_{3}$-washout is absent. [^16] This assumption is conservative but ultimately it is irrelevant whether $N_{3}$-washout is included or not, the predictions on the low energy neutrino parameters change just slightly ST . Because of the RH neutrino inverse processes, the pre-existing asymmetry will undergo a dynamical evolution that has to take into account also the different flavour regimes. Finally, the relic value of the pre-existing asymmetry can be also expressed as the sum of three contributions from each flavour i.e., $N_{B-L}^{\rm p,f}=\sum_{\alpha}\,N_{\Delta_{\alpha}}^{\rm p,f}$. Neglecting flavour coupling, the expressions for the $N_{\Delta_{\alpha}}^{\rm p,f}$ are given by ST , chianese
@@
In this expression the quantities $p^{0}_{{\rm p}\tau}$ and $p^{0}_{{\rm p}\tau_{2}^{\bot}}$ denote the fractions of the initial pre-existing asymmetry in the tauon flavour and in the flavour $\tau_{2}^{\bot}$, where, as already defined, $\tau_{2}^{\bot}$ is the electron and muon flavour superposition component in the leptons produced by $N_{2}$-decays (or equivalently the flavour component that is washed-out in the inverse processes producing $N_{2}$). [^17] Notice that in general $p^{0}_{{\rm p}\tau}+p^{0}_{{\rm p}\tau_{2}^{\bot}}\neq 1$ since in general the projection of the pre-existing asymmetry flavour direction on the $\tau^{\bot}$ plane, is not parallel to $\tau^{\bot}_{2}$. We also introduced the two quantities $p^{0}_{\alpha\tau_{2}^{\bot}}\equiv K_{2\alpha}/(K_{2e}+K_{2\mu})\;(\alpha=e,\mu)$. These give the fractions of the $\alpha$ flavour in the $\tau_{2}^{\bot}$ component of the lepton state $|\ell_{2}\rangle$ produced by $N_{2}$-decays. The terms $\Delta p_{{\rm p}\tau},\Delta p_{{\rm p}e}$ and $\Delta p_{{\rm p}\mu}$ are the so-called phantom terms and such that $\Delta p_{{\rm p}e}+\Delta p_{{\rm p}\mu}+\Delta p_{{\rm p}\tau}=0$. They are phantom terms taking into account the possibility that the mechanism that created the pre-existing asymmetry produced lepton states that are not exactly the CPC\!P conjugated of the anti-lepton states. It can be crossed-checked that in Eq. (53), if one switches-off the washout terms, then the sum of the three final flavoured pre-existing asymmetries is equal to the total initial pre-existing asymmetry, as it should be.
@@
The strong thermal leptogenesis condition requires that the relic value of the pre-existing asymmetry is negligible with respect to the final value of the asymmetry produced by $N_{2}$-decays, for generic initial values of the flavoured pre-existing asymmetries. For definiteness, we will impose a condition $N_{B-L}^{\rm p,f}<0.1\,N_{B-L}^{\rm lep,f}$.
@@
The only possibility to have successful strong thermal leptogenesis is just the $N_{2}$ tauon-dominated leptogenesis scenario problem that is also representing the bulk of solutions in SO10INLEP, as we discussed. This is quite an intriguing coincidence. The reason why only tauon-dominates solutions can realise strong thermal leptogenesis can be understood directly from Eq. (53). One can indeed have in this case $K_{2\tau},K_{1\mu},K_{1e}\gg 1$ and $K_{1\tau}\lesssim 1$ in a way to wash out at $T\sim M_{2}$ the tauonic component of the pre-existing asymmetry and at $T\sim M_{1}$ also the electronic and muonic components. At the same time, having $K_{1\tau}\lesssim 1$, the tauonic asymmetry produced from leptogenesis is the only that survives the $N_{1}$-washout.
@@
There is only a very special caveat chianese . If one imposes $(1-p^{0}_{\mu\tau_{2}^{\bot}})=K_{2e}/(K_{2e}+K_{2\mu})\lesssim 0.1N_{B-L}^{\rm lep,f}/N_{B-L}^{\rm p,i}\lll 1$ and $K_{1\mu}\lesssim 1$, instead of $K_{1\mu}\gg 1$, one can have a strong thermal muon-dominated solution. We noticed that $K_{2e}/(K_{2e}+K_{2\mu})\sim\theta_{\rm c}^{2}\sim 0.05$ is indeed small but in this case, for a meaningfully large value of pre-existing asymmetry $\sim 10^{-3}$, as we will assume, one needs this ratio at least four orders of magnitude smaller. These are incredibly special solutions, unless they can be justified within a particular model. In chianese these solutions were found with a tiny $\sim 10^{-9}$ rate. Therefore, they have to be considered more like oddities, rather than real ST-SO10INLEP solutions.
@@
By imposing a strong thermal leptogenesis condition in addition to successful SO10INLEP, one singles out a special subset of solutions within those we obtained just imposing successful SO10INLEP. We have assumed as benchmark value $N^{\rm p,i}_{B-L}=10^{-3}$. In the scatter plots, in Fig. 1 and Figs. 3–9, these solutions are denoted by blue points. The success rate for these solutions in our scan is $10^{-6}$. Our results fully confirm previous findings ST , decrypting , full , chianese . It is still useful to notice some important aspects:
@@
  • From Fig. 6 one can see how the requested values of the Majorana phases for ST-SO10INLEP become so constrained that ST-SO10INLEP would be a way to determine their values.
  • The absolute neutrino mass scale range is quite restricted. For the benchmark value we assumed, $N^{\rm p,i}_{B-L}=10^{-3}$, one has that successful ST-SO10INLEP requires $m_{1}=(10$30)30)\,meV and, from Fig. 8, one can see a corresponding similar range for $m_{ee}$. It is interesting that both cosmological observations and $0\nu\beta\beta$ experiments are now starting to test this range.
  • The bulk of solutions realising ST-SO10INLEP are of $\tau_{A}$ type. The $\tau_{B}$ type solutions have the problem that $K_{1\mu}$ tends to be too small to washout the muonic component of the pre-existing asymmetry. For example for $V_{L}=I$ one finds $K_{1\mu}\lesssim 4$ decrypting and in this case they would be absolutely unable to yield strong thermal solutions. However, turning on angles in $V_{L}$ values as large as $K_{1\mu}\sim 10$ are possible and for this reason some $\tau_{B}$ strong thermal solutions do exist, though they contribute only $\sim 1\%$ to the total number of solutions.
@@
Figure 10: ST-SO10INLEP allowed region in the plane $\delta$ versus $\theta_{23}$. The red line delimits the region when flavour coupling is neglected, the coloured region is the result in the case when flavour coupling is included. The colour code denotes point density: lighter blue means higher density of solutions. The thin dashed lines are the $1\sigma,2\sigma$ and $3\sigma$ experimental constraints from nufit24 .
The result is shown in Fig. 10 and indicated by a red line. This well reproduces the result found in chianese . The small differences can be ascribed to the updated experimental ranges for neutrino oscillation parameters in Eq. (6) we are adopting. The thin dashed lines are the the $1\sigma,2\sigma$ and $3\sigma$ experimental constraints from nufit24 . One can see how ST-SO10INLEP is in nice agreement with latest experimental results within $1\sigma$. The fact that the current best fit falls in the first octant releases the tension that was found with old experimental data in chianese . The best fit $\delta$ value is actually too low but this is very sensitive to the employed data sets in the analysis, mainly because of an existing tension between T2K and NOVA data. In the left panel of Fig. 11 we show
@@
plots of $\eta_{B}^{\rm lep}$ versus $m_{1}$ for the indicated values of $\theta_{23}$ and for fixed $\delta=-50^{\circ}$, corresponding to the value allowing the maximum value of $\theta_{23}$. All other parameters are fixed to values that maximise the asymmetry and are given in Table 3. One can see how for increasing values of $\theta_{23}$ the peak value of $\eta_{B}^{\rm lep}$ decreases. This is well understood since the asymmetry is approximately $\propto 1/\sin^{4}\theta_{23}$ decrypting , full . One can see how the maximum value of $\theta_{23}$ we obtain is $46^{\circ}$. This is consistent with the value obtained in chianese and it clearly shows how the second octant is incompatible with ST-SO10INLEP.
@@
No F.C. $33.62^{\circ}$ $9.09^{\circ}$ See Figure $49^{\circ}$ 0.312 0.834 $4.11^{\circ}$ $0.183^{\circ}$ $2.41^{\circ}$ 1.92 1.24 0.18
With F.C. $33.62^{\circ}$ $9.09^{\circ}$ See Figure $49^{\circ}$ 0.312 0.834 $4.11^{\circ}$ $0.183^{\circ}$ $2.41^{\circ}$ 1.92 1.24 0.18
@@ 4.2 Including flavour coupling
We have now finally to discuss the impact of flavour coupling on ST-SO10INLEP. This is one of the main motivations of the paper. A simplistic, but legitimate, expectation is that flavour coupling might completely disrupt ST-SO10INLEP. The reason is that a large pre-existing asymmetry, thanks to flavour coupling, can find its way to survive into the tauon flavour where the washout is weak. As we are going to explain, this expectation is not met.
First of all, we have to generalise the expression (53) for the relic value of the pre-existing asymmetry. If we start with an initial pre-existing asymmetry $N^{\rm p,i}_{B-L}$, that we assume to be generated at some temperature above $\sim 10^{12}\,{\rm GeV}$, below $10^{12}\,{\rm GeV}$ the tauon component of lepton states is measured by tauon charged lepton interactions and the pre-existing asymmetry will be the sum of a tauon component, given by $p^{0}_{{\rm p}\tau}\,N^{\rm p,i}_{B-L}$, and a component along a flavour direction given by the projection of the flavour composition of pre-existing leptons on the $e$-$\mu$ flavour plane given by $(1-p^{0}_{{\rm p}\tau})\,N^{\rm p,i}_{B-L}$. In addition, in general, there should be also the phantom terms that however we neglect just to shorten the notation (in any case results are not sensitive to them). Similarly to the solution for the asymmetry generated by leptogenesis, at $T\sim M_{2}$ the $N_{2}$-washout, in the presence of flavour coupling, uncouples in a rotated flavour basis $\tau^{\prime}$ and $\tau^{\bot^{\prime}}_{2}$. The difference is that now there is no source term and, therefore, at the end of the washout processes the washout is described by exponential factors and, rotating back to the standard flavour basis, one finds
@@
where the matrix $U$ is the same we introduced in Eq. (46) when we introduced flavour coupling for the asymmetry produced by $N_{2}$-decays. It can be noticed that, for $U=I$, one recovers the result in the absence of flavour coupling and $N_{1}$-washout in Eq. (53). Also notice that the component of the pre-existing asymmetry along the flavour $\tau_{2^{\bot}}^{\bot}$, orthogonal to $\tau$ and $\tau_{2}^{\bot}$, escapes the washout at $T\sim M_{2}$ Barbieri:1999ma , Engelhard:2006yg .
@@
Using now this much more intricate expression, we have obtained ST-SO10INLEP solutions shown, again with blue points, in the right panels of Fig. 1 and Figs. 2–8. The first thing to notice is that the region survives and in fact it even slightly increases. In particular, the most noticeable change is that the lower bound on $m_{1}$ gets relaxed and from $m_{1}\gtrsim 8\,{\rm meV}$, obtained neglecting flavour coupling, we have now found $m_{1}\gtrsim 3\,{\rm meV}$. One can also see from Fig. 8 that, corresponding, the lower bound on $m_{ee}$ gets relaxed from $m_{ee}\gtrsim 9\,{\rm meV}$ to $m_{ee}\gtrsim 4\,{\rm meV}$. One can also see from Fig. 6 that now the two disconnected subregions get linked by the appearance of the new solutions.
This implies, that in fact the washout of a pre-existing asymmetry, gets even easier when flavour coupling is taken into account, contrarily to naive expectations. The reasons why that happens were already partially envisaged in full . Even though flavour coupling induces part of the pre-existing asymmetry in the electron and muon flavours (that survived the $N_{2}$-washout) to be transferred in the tauon flavour, where the washout is absent ($K_{1\tau}\ll 1$) to let the asymmetry produced from leptogenesis to survive, the off-diagonal term in the V and $V^{-1}$ matrices responsible for the coupling of $e$ ($\mu$) flavour to the $\tau$ flavour is proportional to $K_{1\tau}/K_{1e}(K_{1\tau}/K_{1\mu})$ and so intrinsically small. This means that the same conditions that are necessary to have strong thermal leptogenesis are also sufficient to protect it from the potential disruptive effect of flavour coupling. In addition flavour coupling is such that that the flavoured decay parameters responsible for the washout of the pre-existing electron and muon asymmetries are not the standard $K_{1e}$ and $K_{1\mu}$ but these are replaced by $K_{1e}^{\prime\prime}$ and $K_{1\mu}^{\prime\prime}$. The origin of the lower bound on $m_{1}$ was studied in detail in sophie , decrypting and it was due to the fact that $K_{1e}\simeq m_{ee}/m_{\star}\simeq 0.8\,m_{1}/m_{ee}$. For an initial electronic pre-existing asymmetry $\sim 10^{-4}$, the strong thermal conditions implies $K_{1e}\gtrsim 8$ that translates immediately into $m_{ee}\gtrsim 8\,{\rm meV}$ and $m_{1}\gtrsim 9\,{\rm meV}$. However, when flavour coupling is included the washout of the electron asymmetry is described by $K_{1e}^{\prime\prime}$ that is a linear combination primarily of $K_{1e}$ but with a small contribution from $K_{1\mu}\simeq(m_{\rm sol}s^{2}_{12}+m_{\rm atm}s_{13}c_{23})^{2}/m_{\star}m_{\rm sol}s^{2}_{12}\simeq 25$. This contamination from a large $K_{1\mu}$ makes possible to have smaller values of $K_{1e}$ and, therefore, of $m_{ee}$ and $m_{1}$.